OpenStudy (anonymous):

the area of a rectangular pool by the trinomial 2y^2+y-3. what are the possible dimensions of the pool?

OpenStudy (anonymous):

@praxer

OpenStudy (praxer):

2y^2+3y-2y-3 =2y^2-2y+3y-3 =2y(y-1)+3(y-1) =(2y+3)(y-1)

OpenStudy (anonymous):

answer choices (-y-1) and (-2y-3) y-1) and (2y+3) (2y-1) and (2y+3) (2y+1) and (y-3)

OpenStudy (anonymous):

crazyinlove you look kind of like wwe jojo but anyways here is the link http://openstudy.com/study#/updates/518bcda8e4b062a8d1d95716

OpenStudy (anonymous):

what

OpenStudy (anonymous):

@Tysonkiddandnattie

OpenStudy (anonymous):

thank u i got the answer

OpenStudy (anonymous):

ur welcome